1a.) You receive an infusion of TPN 1.5L to run at 65 mL/hr. How long will it take for the entire bag to infuse?

1b.) Your patient is prescribed Amoxicillin 125 mg/kg/day to be administered in divided doses every 6 hours. The patient weighs 110 lbs. You have available a bottle of Amoxicillin 500mg/10mL. It is 0600 and a dose of the Amoxicillin is scheduled to be administered. How many mL will the patient receive? Round answer to the nearest whole number.

1c.) The prescriber has written an order for the patient to receive Ancef 2 grams in 100 mL of 0.9% NS to infuse over 30 mins. At what rate will you set the IV pump?
Answer #1

ANSWER.

1) a) Calculation of time required to complete the infusion.

          Formula

               Rate (ml/hr) = Total volume(ml) /Time(hr).

           So

                 Time(hr) = Total volume(ml) /Rate (ml/hr).

     Here

              Volume (ml) = 1500 ml (1.5 L)

               Rate(ml/hr) = 65 ml/hr.

   As per above formula

                  Time(hr) = 1500 / 65

                                 = 23.07

                                 = 23 hrs.

1) b) Calculation of volume(ml) of drug the patient receive.

           Formula

               Volume(ml) = Desired dose(D) / Dose in hand (H) * Quantity(Q).

         Here

               Ordered dose = 125 mg/kg/day.

                Weight of the patient = 110 lbs.

        Convert lbs to kg for calculation

                        1 lbs = 0.454 kg.

             Weight of the patient in kg = 110 * 0.454

                                                        = 49.9 kg.

               Desired dose (D) = 125 * 49.9

                                            = 6237.5 mg

               Dose in hand (H) = 500 mg.

                        Quantity (Q) = 10 ml.

   As per above formula

               Volume of drug required = D/H * Q

                                                      = 6237.5 / 500 * 10

                                                      = 12.47 * 10

                                                      = 124.7

                                                      = 124.7 ml per day.

        Volume(ml) of drug received per dose.

               Volume of drug received per day = 125 ml.

               Number of times drug delivered in a day = Day (24 hrs) /Every 6 hrs.

                                                                                = 4

            So

                Volume(ml) of drug received per dose = 124.7 / 4

                                                                                    = 31.2

                                                                                    = 31 ml.

1) c) Calculation of rate of infusion.

                    Rate (ml/hr) = Total volume(ml) /Time(hr).

           Here

                   Total volume(ml) = 100 ml.

                        Time(hr) = 0.5 hr (30 min).

      As per above formula

                  Rate (ml/hr) = 100 / 0.5

                                          = 200 ml/hr.